The Lakeside Company uses a weighted-average process costing system. The following data are available:

Beginning inventory -0-
Units started in production 23,000
Units finished during the period 17,500
Units in process at the end of the period
(complete as to materials,
¼ complete as to labor and overhead)

5,500
Cost of materials used $ 53,750
Labor and overhead costs $ 55,755

Cost per equivalent unit of materials is:

  • $2.95.

  • $2.66.

  • $2.42.

  • $2.34.
Answer #1

equivalent unit of material = 17500+5500 = 23000

Cost per equivalent unit of material = 53750/23000 = 2.34

So answer is d) $2.34